Sulfur trioxide decomposes at high temperature in a sealed container according to the reaction 2SO3(g) ⇆ 2SO2(g) + O2(g)


Initially, the vessel is charged at 1000 K with SO3(g) at a partial pressure of 0.500 atm. At equilibrium the SO3 partial pressure is 0.200 atm. At equilibrium the SO3 partial pressure is 0.200 atm. Calculate the value of Kp at 1000 K.


Strategy:

Using the equilibrium table, tabulate the given intial and equilibrium partial pressures. Denote the changes in partial pressure using a variable as guided by stoichiometric coefficients in the balanced chemical equation. Solve for the unknown to calculate for the equilibrium partial pressures of SO2 and O2 and then solve for Kp.


1
Expert's answer
2022-06-06T17:05:04-0400

0,5-0,2=0,3 reacted

2SO3 = 2SO2 + O2

2 SO3 —> 2 SO2 — 1 O2

0,3 SO3 —> X — Y


X= 0,3*2/2= 0,3 SO2

Y= 0,3*1/2= 0,15 O2


"K_p = \frac{[SO_2]²[O_2]} {[SO_3]²}"


"K_p = \frac{[0,3]²[0,15]} {[0,2]²}= 0,3375"


Answer: Kp = 0,3375